OpenStudy (anonymous):

The area of a rectangular wall in a classroom is 165 square feet. Its length is 4 feet longer than its width. Find the width of the wall of the classroom. This question is in "Completing the Square" section

OpenStudy (anonymous):

Well, we could use the W as X and L = X + 4

OpenStudy (anonymous):

x(x + 4) = 165 {area of a rectangle is width x length} x² + 4x = 165 {used distributive property} x² + 4x - 165 = 0 {subtracted 165 from both sides} (x + 15)(x - 11) = 0 {factored into two binomials} x + 15 = 0 or x - 11 = 0 {set each factor equal to 0} x = -15 or x = 11 {subtracted 15 and added 11} x = 11 {width of a rectangle cannot be negative}

OpenStudy (anonymous):

Hope it helped. ^^